A right triangle has a perimeter of 3 meters. It's hypotonuse is 130 cm. What are the lengths of the other sides?
Answer by fractalier(6550)

Let the two legs be a and b.
So we have a + b + 130 = 300 (300 cm is 3 m) and from the Pythagorean Theorem, we have
a^2 + b^2 = 130^2
Now solve the first equation for a and plug it in to the second equation, and solve using the factoring technique, so
a + b = 170
a = 170 - b
(170 - b)^2 + b^2 = 16900
28900 - 340b + b^2 + b^2 = 16900
2b^2 - 340b + 12000 = 0
b^2 - 170b + 6000 = 0
(b - 120)(b - 50) = 0
b = 120 or b = 50 thus
a = 50 or a = 120